# Scattering of non-relativistic finite-size particles and puffy dark matter direct detection

### v1

@@ 3.3 Detection of nugget-type dark matter
The finite-size effect of particles can significantly impact their scattering cross section. For the scattering of large-size particles, as discussed in the preceding section, the cross section lies within the Born approximation regime. This implies that the interaction is effectively confined within the internal structure of the particles, resembling a short-range force. In this regime, the potential energy is much smaller than the kinetic energy, and thus the quantum mechanical and quantum field treatments are effectively equivalent. In contrast, for the scattering of small-size particles, this equivalence no longer holds, and accurate cross sections must be computed using partial wave analysis. In the following we focus on the scattering between a finite-size bound state of dark matter composed of a small number of constituent particles, such as a nugget-like dark matter with a small $N$, with a nucleus Wise:2014ola ; Wise:2014jva ; Gresham:2017cvl ; Bai:2018dxf . Ref. Gresham:2017zqi utilizes relativistic mean field theory and gives a systematic computation of nugget properties. For nugget dark matter with a small $N$, the constituent density is low and the constituents are non-relativistic. In this case, the self-interaction force range between dark matter particles is large (with the dark matter particle radius smaller than the force range, $R_{\chi}\leq m_{\phi}^{-1}$). As a result, the effect of the mediator particle mass is not significant. The mass of the dark matter particle is approximately $Nm_{\chi}$. Using the non-relativistic formula for a fermi gas, the radius of the dark matter particle can be estimated as $R_{\chi}\simeq[81\pi^{2}/(4Ng_{\rm dof}^{2}\alpha_{\chi}^{3}m_{\chi}^{3})]^{1/3}$ (see Ref. Wise:2014jva for details). Here, $\alpha_{\chi}$ is the coupling constant for the self-interaction of the dark matter particles, $g_{\rm dof}$ is the number of degrees of freedom for the fermion field. The stability condition of the bound state provides a relation between the mass and radius of the nugget dark matter particle. Furthermore, in the detection of nugget-type dark matter, the use of partial wave analysis allows for a more predictive parameter space for dark matter–nucleus scattering.
@@
Next, we consider a bound-state dark matter particle with constituent number $N=5$ or $N=10$. Given the self-coupling constant $\alpha_{\chi}=\alpha=0.01(0.1)$, we substitute the potential function Eq.(25) into the Schrödinger equation Eq.(21). By scanning the parameter space $(m_{\phi},m_{\chi})$, we obtain the nugget dark matter–xenon nucleus scattering cross section. As shown in Fig. 8, the nugget dark matter–xenon nucleus scattering cross section does not exceed $10^{-16}\rm cm^{2}$ for $N=5$ and $N=10$. The similar contour shapes in the left and right panels indicate that a small number of bound-state constituents has little effect on the nugget dark matter–nucleus scattering cross section. Moreover, the larger the coupling constant, the larger the allowed dark matter parameter space. For $\alpha=0.01(0.1)$, the stability condition of the nugget-type bound state requires the dark matter particle mass to be greater than $0.1(0.01)\rm GeV$. In addition, one can also observe the appearance of resonance structures in the cross section from both panels.
@@ 4 Conclusion
Unlike the approach based on non-relativistic effective operators, we directly constructed the interaction potential between finite-size particles in coordinate space and then solved the Schrödinger equation to obtain the scattering cross section between non-relativistic finite-size particles. We found that due to the finite size effect of particles, the interaction potential between finite-size particles no longer exhibits the singular behavior of the point-particle Yukawa potential; instead, it approaches a constant value at the origin. The interaction potential between particles of unequal sizes was found to have a more complex dynamics. Applying this method to the study of dark matter direct detection, we obtained the following observations: (i) Even for point-like dark matter particles, the finite size of the target nucleus leads to a dark matter–nucleus scattering cross sections which can be classified into Born, resonant, and classical regimes; (ii) For the scattering between a large-size dark matter particle and a nucleon, the cross section lies entirely within the Born regime. Since in this case the potential energy between particles is much smaller than the kinetic energy, the results agree with those obtained from quantum field theory at the tree level; (iii) For a small-size puffy dark matter particle, the complex dynamics of the interaction potential gives rise to rich phenomenology in the dark matter–nucleus scattering cross section; (iv) For the direct detection of nugget-type dark matter with a small number of constituent particles, the stability condition of the dark matter bound-state imposes constraints that define the viable parameter space and the corresponding dark matter–nucleus scattering cross section region.

### v2

@@ 3.3 Detection of nugget-type dark matter
The finite-size effect of particles can significantly impact their scattering cross section. For the scattering of large-size particles, as discussed in the preceding section, the cross section lies within the Born approximation regime. This implies that the interaction is effectively confined within the internal structure of the particles, resembling a short-range force. In this regime, the potential energy is much smaller than the kinetic energy, and thus the quantum mechanical and quantum field treatments are effectively equivalent. In contrast, for the scattering of small-size particles, this equivalence no longer holds, and accurate cross sections must be computed using partial wave analysis. Beyond that, these results can also be seen in Figs. 5-7. The scattering cross sections between small-size particles occur in the resonance and classical regions. As for the Bonn region, it corresponds to the case of large-size superheavy dark matter particles studied in the literature using quantum field theory to calculate the scattering cross sections Cappiello et al. (2021); Boukhtouchen et al. (2025); Clark et al. (2020); Carney and others (2023); Aggarwal and Raj (2025); Aalbers and others (2024). Similarly, current research on nugget type dark matter focuses on the supermassive mass range Wise and Zhang (2015, 2014); Gresham et al. (2018); Bai et al. (2019). Our work, focusing on the direct detection of bound state dark matter composed of a small number of particles, is a supplement to the direct detection research of nugget type dark matter. In the following we focus on the scattering between a finite-size bound state dark matter particle composed of a small number of constituent particles, such as a nugget-type dark matter particle with a small $N$, and a finite-size nucleus. Ref. Gresham et al. (2017) utilizes relativistic mean field theory and gives a systematic computation of nugget properties. For nugget dark matter with a small $N$, the constituent density is low and the constituents are non-relativistic. In this case, the self-interaction force range between dark matter particles is large (with the dark matter particle radius smaller than the force range, $R_{\chi}\leq m_{\phi}^{-1}$). As a result, the effect of the mediator particle mass is not significant. The mass of the dark matter particle is approximately $Nm_{\chi}$. Using the non-relativistic formula for a fermi gas, the radius of the dark matter particle can be estimated as $R_{\chi}\simeq[81\pi^{2}/(4Ng_{\rm dof}^{2}\alpha_{\chi}^{3}m_{\chi}^{3})]^{1/3}$ (see Ref. Wise and Zhang (2014) for details). Here, $\alpha_{\chi}$ is the coupling constant for the self-interaction of the dark matter particles, $g_{\rm dof}$ is the number of degrees of freedom for the fermion field. The stability condition of the bound state provides a relation between the mass and radius of the nugget dark matter particle. Furthermore, in the detection of nugget-type dark matter, the use of partial wave analysis allows for a more predictive parameter space for dark matter–nucleus scattering.
@@
Next, we consider a bound-state dark matter particle with constituent number $N=5$ or $N=10$. Given the self-coupling constant $\alpha_{\chi}=\alpha=0.01(0.1)$, we substitute the potential function Eq.(25) into the Schrödinger equation Eq.(21). By scanning the parameter space $(m_{\phi},m_{\chi})$, we obtain the nugget dark matter–xenon nucleus scattering cross section. As shown in Fig. 8, the nugget dark matter–xenon nucleus scattering cross section does not exceed $10^{-16}\rm cm^{2}$ for $N=5$ and $N=10$. The similar contour shapes in the left and right panels indicate that a small number of bound-state constituents has little effect on the nugget dark matter–nucleus scattering cross section. Moreover, the larger the coupling constant, the larger the allowed dark matter parameter space. For $\alpha=0.01(0.1)$, the stability condition of the nugget-type bound state requires the dark matter particle mass to be greater than $0.1(0.01)~\rm GeV$. In addition, one can also observe the appearance of resonance structures in the cross section from both panels.
@@ 4 Conclusion
Unlike the approach based on non-relativistic effective operators, we directly constructed the interaction potential between finite-size particles in coordinate space and then solved the Schrödinger equation to obtain the scattering cross section between non-relativistic finite-size particles. We found that due to the finite size effect of particles, the interaction potential between finite-size particles no longer exhibits the singular behavior of the point-particle Yukawa potential; instead, it approaches a constant value at the origin. The interaction potential between particles of unequal sizes was found to have a more complex dynamics. Applying this method to the study of dark matter direct detection, we obtained the following results: (i) For a point-like dark matter particle and a finite-size target nucleus, our re-calculation results agree with Ref. Xu and Farrar (2023), which shows the dark matter–nucleus scattering cross section can be classified into Born, resonant, and classical regimes; (ii) For the scattering between a puffy dark matter particle with a large size-to-range ratio and a finite-size nucleus, the cross section lies entirely within the Born regime. Since in this case the potential energy between the particles is much smaller than the kinetic energy, the results agree with those obtained from quantum field theory at the tree level; (iii) For a puffy dark matter particle with a small size-to-range ratio, the complex dynamics of the interaction potential gives rise to rich phenomenology in the scattering cross section between dark matter and finite-size nucleus, which can also be classified into Born, resonant, and classical regimes; (iv) For the direct detection of nugget-type dark matter with a small number of constituent particles, due to the stability condition of the dark matter bound-state, the viable dark matter–nucleus scattering cross section region is specified.